# Batch Predictive Inference

### v4

@@ 2.4 Inference under covariate shift
In this section, we develop methods for batch predictive inference under covariate shift. This refers to the following distribution of the data points:
@@
Unfortunately, it is known that even in the case of a single test point, constructing a completely distribution-free prediction set under covariate shift is not possible unless the prediction sets are uninformative (cover each fixed value with probability $1-\alpha$)  (Qiu et al., 2023; Yang et al., 2023+). Therefore, as for the case of standard conformal prediction  (Tibshirani et al., 2019), we consider the setting of a known likelihood ratio $dP/dQ$. This setting is of broad interest, arising in randomized trials (e.g., Friedman et al., 2010; Armitage et al., 2013, etc.). and two-phase sampling studies (e.g., Hansen and Hurwitz, 1946; Shrout and Newman, 1989, etc.). As in other problems in conformal prediction, studying the setting where the likelihood ratio is unknown and needs to be estimated might require significant additional development, as well as a different set of tools (Qiu et al., 2023; Yang et al., 2023+).
2.4.1 Reformulation as a missing data problem
To enable a concise argument, it helps to reformulate the problem as a missing data problem. Let $A\in\{0,1\}$ be the binary variable that indicates whether or not the outcome $Y$ is observed. Then the set of all observed data points $(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n}),X_{n+1},X_{n+2},\ldots,X_{n+m}$ can equivalently be viewed as having $n+m$ tuples $(X_{i},A_{i},Y_{i}A_{i})_{1\leqslant i\leqslant n+m}$. The feature distributions $P_{X}$ and $Q_{X}$ in (17) correspond to the conditional distributions $P_{X\mid A=1}$ and $P_{X\mid A=0}$, respectively. Thus, we can rewrite the model (17) as
and the target coverage guarantee (18) can be written as
A simple approach one could consider is to extend weighted split conformal prediction. However, as we show in Section A.4, this approach suffers from a similar issue as the standard extension of split conformal prediction. Unless $n\gg m$, it typically results in large prediction sets that can cover the entire range of the random variable of interest.
2.4.2 Proposed method: batch PI with rejection sampling
Algorithm 7 Batch Predictive Inference under Covariate Shift
As an alternative approach, we consider constructing an exchangeable dataset via rejection sampling, as it has been done for standard conformal prediction in Park et al. (2022a); Qiu et al. (2023), and then applying the batch PI procedure.
Suppose we have access to the conditional distribution $P_{A\mid X}$ (again, for some possibly unknown value of $\mathbb{P}\left\{{A=1}\right\}$). We draw a subset of the calibration data set as follows. For each $i=1,2,\ldots,n$, draw
The Bernoulli distribution described above is well-defined for any value of $X_{i}$ if $p_{A\mid X}(x)>0$ for all $x\in\mathcal{X}$. In fact, for our coverage result, we will require a slightly stronger condition:
Condition 2.
This sampling scheme was previously discussed in Park et al. (2022a), and intuitively, it constructs a subset of the calibration set that mimics the distribution of the test set through reweighting based on the propensity score. Let $\widetilde{\mathcal{D}}_{n}$ be the subset of the calibration data defined as
Corollary 3.
Under Conditions 1 and 2, with $\widetilde{\mathcal{D}}_{n}$ constructed by (21), the batch PI prediction set $\widehat{C}^{\textnormal{bPI-CovShift}}(\mathcal{D}_{n}):=\widehat{C}(\widetilde{\mathcal{D}}_{n})$ based on (8) satisfies $\mathbb{P}\left\{{g(\{Z_{n+1},\ldots,Z_{n+m}\})\in\widehat{C}(\widetilde{\mathcal{D}}_{n})}\ \middle|\ {A_{1:n},B_{1:n}}\right\}\geqslant 1-\alpha$, where the probability is taken with respect to the model (19).
Similarly, we can combine rejection sampling and the procedure (6) to conduct inference on multiple quantiles of test scores under covariate shift. In general, rejection sampling translates any procedure designed for i.i.d. data to a procedure suitable for data with covariate shift. The procedure $\widehat{C}(\widetilde{\mathcal{D}}_{n})$ is an application of this approach to batch PI. Since rejection sampling reduces the sample size, using naive procedures such as split conformal prediction may yield uninformative prediction sets after rejection sampling, even if the original calibration set is large. The batch PI procedure addresses this issue as its usefulness does not depend heavily on the ratio of calibration to test sizes.
@@ 3 Use cases
In this section, we discuss use cases of batch PI: (1) simultaneous predictive inference with dataset-conditional coverage; (2) selection of individuals with error control –—both based on inference on one quantile; and (3) inference on counterfactual variables. All three will be illustrated empirically in Section 4.
@@ 3.1 Simultaneous predictive inference of multiple unobserved responses
Consider constructing prediction sets $\widehat{C}_{n}(X_{n+1}),\widehat{C}_{n}(X_{n+2}),\ldots,\widehat{C}_{n}(X_{n+m})$ for $Y_{n+1},Y_{n+2},\ldots,Y_{n+m}$ respectively, such that most of the unobserved outcomes are covered by their corresponding prediction sets. A simple approach is to construct standard split conformal prediction sets, leading to marginal coverage for each prediction set, i.e., $\mathbb{P}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}\geqslant 1-\alpha,\text{ for all }j=1,2,\ldots,m$.
However, this does not characterize the simultaneous—joint—behavior of the prediction sets. For instance, it does not directly guarantee how many of the test outcomes will be covered. Since each marginal coverage guarantee is with respect to the distribution of $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+j},Y_{n+j})$, the $m$ coverage events $\{\{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})\},j=1,2,\ldots,m\}$ have a joint distribution with a potentially complex dependence structure.
The distribution of the coverage $\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}$ was determined by Marques F. (2025); Huang et al. (2024), and this enables constructing prediction sets various guarantees. Our goal is to construct prediction sets with the following probably approximately correct (PAC)-type (Park et al., 2020) guarantees:
$\mathbb{P}\left\{{\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}\geqslant 1-\delta}\right\}\geqslant 1-\alpha,$ (22)
@@
$\widehat{C}_{n}(x)=\left\{y\in\mathcal{Y}:s(x,y)\leqslant Q_{1-\alpha}\left(\sum_{k=1}^{n+1}\frac{\binom{k+m_{\delta}-2}{m_{\delta}-1}\binom{n+m-k-m_{\delta}+1}{m-m_{\delta}}}{\binom{n+m}{m}}\cdot\delta_{k}\right)\right\}.$ (23)
As a consequence of our general results, we find that if $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),$ $(X_{n+1},Y_{n+1}),$ $\ldots$,… , $(X_{n+m},Y_{n+m})$ are exchangeable, then the prediction set $\widehat{C}_{n}$ from (23) satisfies the guarantee (22). This follows directly by applying the result in Section 2.2.1.
Remark 3 (Comparison with the PAC guarantee for calibration-conditional coverage).
Consider the setting where the data points are i.i.d. Let $C_{j}={\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}$ denote the coverage indicator for the $j$th test point, $j\in[m]$, and let $\mathcal{D}_{\text{cal}}$ denote the calibration set. Then we have $C_{1},\cdots,C_{m}\mid\mathcal{D}_{\text{cal}}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}\textnormal{Bernoulli}(p_{C})$, where $p_{C}=\mathbb{P}\left\{{Y\in\widehat{C}_{n}(X)}\ \middle|\ {\mathcal{D}_{\text{cal}}}\right\}$, and thus $\bar{C}=\frac{1}{m}\sum_{j=1}^{m}C_{j}$ converges to $p_{C}$ almost surely, conditional on $\mathcal{D}_{\text{cal}}$. It follows that
@@
by applying the dominated convergence theorem twice. Therefore, as $m\to\infty$, the prediction set (23) can also be viewed as achieving the PAC guarantee for the calibration conditional-coverage property (Vovk, 2013b; Park et al., 2020) $\mathbb{P}\left\{{p_{C}\geqslant 1-\delta}\right\}\geqslant 1-\alpha$. The advantage of the prediction set (23) is that it controls the coverage rate also for small test sizes $m$.
@@
This problem was also studied previously in Gazin et al. (2024), where the authors further aim to provide uniform control over the false coverage rate. This can be expressed in our notation as:
$\mathbb{P}\left\{{\forall\,\alpha\in(0,1),\;\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}^{(\alpha)}(X_{n+j})}\right\}\geqslant 1-\gamma_{\alpha,\delta}}\right\}\geqslant 1-\delta,$
where $(\gamma_{\alpha,\delta})_{0<\alpha<1}$ is a family of (random) bounds. They provide a concentration inequality-based approach to achieve this stronger notion of coverage, with a score of the form $S_{i}=|Y_{i}-\hat{\mu}(X_{i},\mathcal{D}_{\text{train}},X_{1:n+m})|$—i.e., the score is constructed using the training data, as well as the calibration and test covariates. They also briefly mention the weaker target (equivalent to (22)) in the appendix, providing method based on an implicit formula. It turns out that explicitly restating their approach is equivalent to the method in (23). In other words, applying the batch PI methodology to false coverage control recovers the approach of Gazin et al. (2024) in a simpler explicit form.
3.2 Selection of test individuals
@@
If $\eta=0$, then (25) is equivalent to making at least one false claims with probability at most $\alpha$, $\mathbb{P}\left\{{\sum_{j=1}^{m}{\mathbbm{1}}\left\{{\hat{\mu}(X_{n+j})>\hat{T},Y_{n+j}\leqslant c}\right\}>0}\right\}$; which is analogous to the control of the family-wise error rate (FWER) in multiple hypothesis testing. More generally, (25) is analogous to the control of the $k$-family-wise error rate ($k$-FWER) (Lehmann and Romano, 2005) in multiple hypothesis testing.
@@
3.3 Inference on counterfactual variables
Finally, we consider a randomized trial setting where the underlying data structure is $(X_{i},A_{i},Y_{i}^{a=0},Y_{i}^{a=1})_{1\leqslant i\leqslant n}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P_{X}\times P_{A\mid X}\times P_{Y^{a=0}\mid X}\times P_{Y^{a=1}\mid X}$, where $X$ denotes the feature, $A\in\{0,1\}$ denotes the treatment, and $Y^{a=0}$ and $Y^{a=1}$ denote the counterfactual outcomes under $A=0$ and $A=1$, respectively. We only observe $(X_{i},A_{i},Y_{i})_{1\leqslant i\leqslant n}$, where we assume the consistency condition $Y_{i}=(1-A_{i})Y_{i}^{a=0}+A_{i}Y_{i}^{a=1}$.
We consider the task of inference on the counterfactual outcomes $\{Y_{i}^{a=0}:A_{i}=1\}$ in the treated group. Under the consistency assumption $Y=(1-A)Y^{a=0}+AY^{a=1}$, the problem is equivalent to inference on missing outcomes/test points under covariate shift, with data points $(X_{i},A_{i},A_{i}Y_{i}^{a=0})$. Indeed, we can regard $\{(X_{i},Y_{i}^{a=0}):A_{i}=0\}$ as the calibration set and $\{X_{i}:A_{i}=1\}$ as the test inputs.
Therefore, based on the discussion in Section 2.4, e.g., by applying the procedure (7), we obtain procedures for the following tasks:
  1. 1.
    Inference on the mean of counterfactuals: Construct $\widehat{C}(\mathcal{D}_{n})$ such that $\mathbb{P}\left\{{\frac{1}{N^{1}}\sum_{i:A_{i}=1}Y_{i}^{a=0}\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$, where $N^{1}=|\{i:A_{i}=1\}|$.
  2. 2.
    Inference on the median of counterfactuals: Construct $\widehat{C}(\mathcal{D}_{n})$ such that $\mathbb{P}\left\{{\textnormal{Median}(\{Y_{i}^{a=0}:A_{i}=1\})\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$.
  3. 3.
    Inference on multiple quantiles of counterfactuals: Construct $L,U\in\mathbb{R}^{l}$ such that $\mathbb{P}\left\{{L\preceq(Y_{(\zeta_{1})}^{a=0},\cdots,Y_{(\zeta_{l})}^{a=0})\preceq U}\right\}\geqslant 1-\alpha$, where $Y_{(\zeta)}^{a=0}$ denotes the $\zeta$-th smallest value of $\{Y_{i}^{a=0}:A_{i}=1\}$.

### v5

@@ 2.4 Inference under covariate shift
Here, we develop methods for batch predictive inference under covariate shift. This refers to the following distribution of the data points:
@@
We consider the setting of a known likelihood ratio $dP/dQ$, which is required for nontrivial distribution-free prediction sets even in the setting of one test datapoint (Qiu et al., 2023; Yang et al., 2023+). We develop a method leveraging rejection sampling to obtain an exchangeable dataset, and then applying the batch PI procedure; similarly to Park et al. (2022a); Qiu et al. (2023) for standard conformal prediction. We present more details in Appendix G.
@@ 3 Use cases
In this section, we discuss use cases of batch PI: (1) inference on counterfactual variables; (2) simultaneous predictive inference with PAC-coverage; and (3) selection of individuals with error control–—the latter two are based on inference on one quantile. All three will be illustrated empirically in Section 4.
3.1 Inference on counterfactual variables
We consider a randomized trial setting where the underlying data structure is
$(X_{i},A_{i},Y_{i}^{a=0},Y_{i}^{a=1})_{1\leqslant i\leqslant n}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P_{X}\times P_{A\mid X}\times P_{Y^{a=0}\mid X}\times P_{Y^{a=1}\mid X},$
where $X$ denotes the feature, $A\in\{0,1\}$ denotes the treatment, and $Y^{a=0}$ and $Y^{a=1}$ denote the counterfactual outcomes under $A=0$ and $A=1$, respectively. We only observe $(X_{i},A_{i},Y_{i})_{1\leqslant i\leqslant n}$, where we assume the consistency condition $Y_{i}=(1-A_{i})Y_{i}^{a=0}+A_{i}Y_{i}^{a=1}$.
We consider the task of inference on the counterfactual outcomes $\{Y_{i}^{a=0}:A_{i}=1\}$ in the treated group. Under the consistency assumption, the problem is equivalent to inference on missing outcomes/test points under covariate shift, with $\{(X_{i},Y_{i}^{a=0}):A_{i}=0\}$ as the calibration set and $\{X_{i}:A_{i}=1\}$ as the test inputs.
Therefore, based on the discussion in Section 2.4, we obtain procedures for the following tasks:
  1. 1.
    Inference on the mean of counterfactuals: Construct $\widehat{C}(\mathcal{D}_{n})$ such that
    $\mathbb{P}\left\{{\frac{1}{N^{1}}\sum_{i:A_{i}=1}Y_{i}^{a=0}\in\widehat{C}(\mathcal{D}_{n})}\right\}$ $\geqslant 1-\alpha$, where $N^{1}=|\{i:A_{i}=1\}|$.
  2. 2.
    Inference on the median of counterfactuals: Construct $\widehat{C}(\mathcal{D}_{n})$ such that
    $\mathbb{P}\left\{{\textnormal{Median}(\{Y_{i}^{a=0}:A_{i}=1\})\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$ [^11] For inference on the median, and more generally on a quantile, we also obtain an upper bound on the coverage based on Corollary 1.
  3. 3.
    Inference on multiple quantiles of counterfactuals: Construct $L,U\in\mathbb{R}^{l}$ such that
    $\mathbb{P}\left\{{L\preceq(Y_{(\zeta_{1})}^{a=0},\ldots,Y_{(\zeta_{l})}^{a=0})\preceq U}\right\}\geqslant 1-\alpha$, where $Y_{(\zeta)}^{a=0}$ is the $\zeta$-th smallest value of $\{Y_{i}^{a=0}:A_{i}=1\}$.
@@ 3.2 Simultaneous predictive inference of multiple unobserved responses
Consider constructing prediction sets $\widehat{C}_{n}(X_{n+1}),\widehat{C}_{n}(X_{n+2}),\ldots,\widehat{C}_{n}(X_{n+m})$ for $Y_{n+1},Y_{n+2},\ldots,Y_{n+m}$ respectively, such that most of the unobserved outcomes are covered by their corresponding prediction sets. A simple approach is to construct standard split conformal prediction sets, leading to marginal coverage for each prediction set, i.e., $\mathbb{P}\{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})\}\geqslant 1-\alpha,\text{ for all }j\in[m]$.
However, this does not characterize the simultaneous—joint—behavior of the prediction sets. For instance, it does not directly guarantee how many of the test outcomes will be covered. Since each marginal coverage guarantee is with respect to the distribution of $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+j},Y_{n+j})$, the $m$ coverage events $\{\{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})\},j\in[m]\}$ have a joint distribution with a potentially complex dependence structure. Nonetheless, the distribution of the coverage $\frac{1}{m}\sum_{j=1}^{m}\mathbbm{1}\{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})\}$ was discussed in Marques F. (2025); Huang et al. (2024), and this enables constructing prediction sets with various guarantees. Our goal is to construct prediction sets with the following probably approximately correct (PAC)-type [^12] This can also be viewed as an analogue of the family-wise error rate, and more generally of the $k$-family-wise error rate from multiple hypothesis testing (Lehmann and Romano, 2005a). For a positive integer $k$, set $\delta=k/m$, so that $k=\delta m$. Then this guarantee is equivalent to $\mathbb{P}\left\{{\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\notin\widehat{C}_{n}(X_{n+j})}\right\}\geqslant k}\right\}\leqslant\alpha$. Now, since $\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\notin\widehat{C}_{n}(X_{n+j})}\right\}$ is the number of errors, this can be viewed as a direct analogue of the $k$-family-wise error rate (Lehmann and Romano, 2005a). In particular, if $k=1$ (i.e., $\delta=1/m$), it can be viewed as an analogue of the family-wise error rate. (Park et al., 2020) guarantees:
$\mathbb{P}\left\{{\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}\geqslant 1-\delta}\right\}\geqslant 1-\alpha,$
@@
$\widehat{C}_{n}(x)=\left\{y\in\mathcal{Y}:s(x,y)\leqslant S_{(r_{\delta,\alpha})}\right\},\text{ where }r_{\delta,\alpha}=Q_{1-\alpha}\bigg{(}\sum_{k=1}^{n+1}\tfrac{\binom{k+m_{\delta}-2}{m_{\delta}-1}\binom{n+m-k-m_{\delta}+1}{m-m_{\delta}}}{\binom{n+m}{m}}\cdot\delta_{k}\bigg{)}.$ (19)
As a consequence of Corollary 1, we establish the following guarantee for the procedure described above.
Corollary 2.
If $(X_{1},Y_{1}),\ldots,(X_{n+m},Y_{n+m})$ are exchangeable, then the prediction set $\widehat{C}_{n}$ from (19) satisfies $1-\alpha\leqslant\mathbb{P}\left\{{\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}\geqslant 1-\delta}\right\}$ $\leqslant 1-\alpha+\varepsilon_{n,m,m_{\delta}},$ where the upper bound holds under the assumption that all the scores $(s(X_{i},X_{i}))_{i\in[n+m]}$ are almost surely distinct, and $\varepsilon_{n,m,m_{\delta}}$ is defined in Corollary 1.
Remark 3 (Comparison with Markov inequality-based approach).
The PAC-type guarantee (18) can also be achieved by applying standard split conformal prediction to each test points separately, at an adjusted level $\delta\cdot\alpha$—i.e., the procedure $\widehat{C}_{n}^{\text{Markov}}=\{y\in\mathcal{Y}:s(x,y)\leqslant S_{(\lceil(1-\delta\alpha)(n+1)\rceil)}\}$. To see this, denote $C_{j}=\mathbbm{1}\{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})\}$. Then, by Markov’s inequality, we have
$\mathbb{P}\bigg{\{}\frac{1}{m}\sum_{j=1}^{m}C_{j}<1-\delta\bigg{\}}=\mathbb{P}\bigg{\{}\frac{1}{m}\sum_{j=1}^{m}(1-C_{j})>\delta\bigg{\}}\leqslant\frac{1}{\delta}\cdot\mathbb{E}\bigg{[}\frac{1}{m}\sum_{j=1}^{m}(1-C_{j})\bigg{]}\leqslant\alpha,$
provided that $\mathbb{E}\left[{C_{j}}\right]\geqslant 1-\delta\alpha$ holds for all $j\in[m]$.
However, this method does not yield a tight bound. In fact, from Proposition 2, it follows that the batch PI-based prediction set $\widehat{C}_{n}(x)$ in (19) is always a subset of $\widehat{C}_{n}^{\text{Markov}}(x)$, for any $x\in\mathcal{X}$. We also provide relevant experimental results in Section 4.1, where we show that our method outperforms the Markov-adjustment based method.
Remark 4 (Comparison with the PAC guarantee for calibration-dataset-conditional coverage).
Consider the setting where the data points are i.i.d. Let $C_{j}={\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}$ denote the coverage indicator for the $j$th test point, $j\in[m]$, and let $\mathcal{D}_{\text{cal}}$ denote the calibration set. Then we have
$C_{1},\cdots,C_{m}\mid\mathcal{D}_{\text{cal}}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}\textnormal{Bernoulli}(p_{C}),\text{ where }p_{C}=\mathbb{P}\left\{{Y\in\widehat{C}_{n}(X)}\ \middle|\ {\mathcal{D}_{\text{cal}}}\right\},$
and thus $\bar{C}=\frac{1}{m}\sum_{j=1}^{m}C_{j}$ converges to $p_{C}$ almost surely as $m\to\infty$, conditional on $\mathcal{D}_{\text{cal}}$. It follows that
@@
by applying the dominated convergence theorem twice. Therefore, as $m\to\infty$, the prediction set (19) converges to achieving the PAC guarantee for the calibration conditional-coverage property (Vovk, 2013b; Park et al., 2020) $\mathbb{P}\left\{{p_{C}\geqslant 1-\delta}\right\}\geqslant 1-\alpha$. The advantage of the prediction set (19) is that it controls the coverage rate also for small test sizes $m$.
@@
This problem was also studied previously in Gazin et al. (2024), where the authors further aim to provide uniform control over the false coverage rate. This can be expressed in our notation as: $\mathbb{P}\left\{{\forall\,\alpha\in(0,1),\;\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}^{(\alpha)}(X_{n+j})}\right\}\geqslant 1-\gamma_{\alpha,\delta}}\right\}$ $\geqslant 1-\delta,$ where $(\gamma_{\alpha,\delta})_{0<\alpha<1}$ is a family of (random) bounds. They provide a concentration inequality-based approach to achieve this stronger notion of coverage, with a score of the form $S_{i}=|Y_{i}-\hat{\mu}(X_{i},\mathcal{D}_{\text{train}},X_{1:n+m})|$—i.e., the score is constructed using the training data, as well as the calibration and test covariates. They also briefly mention the weaker target (equivalent to (18)) in the appendix, providing a method based on an implicit formula—which turns out to be equivalent to the method in (19) after reorganization.
3.3 Selection of test datapoints
@@
If $\eta=0$, then (21) is equivalent to controlling the probablity of making at least one false claims with probability at most $\alpha$, which is analogous to the control of the family-wise error rate (FWER) in multiple hypothesis testing. More generally, (21) is analogous to the control of the $k$-family-wise error rate ($k$-FWER) (Lehmann and Romano, 2005a) in multiple hypothesis testing.
@@
3.3.1 Comparison with p-value-based methods
For the selection problem with the guarantee (21), one might consider first constructing p-values and then applying a standard multiple testing procedure that controls the $k$-FWER (Lehmann and Romano, 2005a). Specifically, we prove the following (see Appendix I for the proof):
Proposition 3.
For the events $E_{1},\ldots,E_{m}$, suppose there exist random variables $p_{1},\ldots,p_{m}$ such that $\mathbb{P}\left\{{p_{j}\leqslant\alpha\text{ and }E_{j}\text{ holds}}\right\}\leqslant\alpha$ for all $\alpha\in(0,1)$ and for all $j\in[m]$. Then the selection rule that selects $E_{j}$ such that $p_{j}\leqslant\frac{(k+1)\alpha}{m}$ controls the $k$-FWER at level $\alpha$, i.e.,
$\mathbb{P}\left\{{\sum_{j=1}^{m}{\mathbbm{1}}\left\{{p_{j}\leqslant\frac{(k+1)\alpha}{m},Y_{n+j}\leqslant c}\right\}\leqslant k}\right\}\geqslant 1-\alpha.$ (22)
The proof is deferred to the Appendix. Note that when $k=0$, the procedure reduces to the simple Bonferroni method, which enjoys FWER control. As the choice of the random variable $p_{j}$, Jin and Candès (2023b) proposes to use the following conformal p-value:
$p_{j}=\frac{\sum_{i=1}^{n}{\mathbbm{1}}\left\{{c-\hat{\mu}(X_{n+j})>Y_{i}-\hat{\mu}(X_{i})}\right\}+1}{n+1}.$ (23)
Jin and Ren (2024) introduces a more powerful conformal p-value, defined as [^13] Jin and Candès (2023b) and Jin and Ren (2024) discuss a more general form of these conformal p-values, of which the p-values (23) and (24) are special cases.
$p_{j}=\frac{\sum_{i=1}^{n}{\mathbbm{1}}\left\{{\hat{\mu}(X_{n+j})<\hat{\mu}(X_{i}),Y_{i}\leqslant c}\right\}+1}{n+1}.$ (24)
However, the multiple testing procedure of Lehmann and Romano (2005a), can be conservative when combined with these p-values. We provide a comparison between these methods and the batch PI-based method through experiments in Section 4.2.
Remark 6.
In the proof of Corollary 3, we show that for a rejection threshold $\hat{T}$ and the rejection rule $\hat{\mu}(X)>\hat{T}$, the $\eta$-FWER is equal to the probability $\mathbb{P}\left\{{S_{(m-\eta)}^{\text{test}}\leqslant\hat{T}}\right\}$. The batch PI procedure finds the optimal threshold $\hat{T}$ based on the exact distribution of the rank of $S_{(m-\eta)}$, and thus the resulting selection rule dominates any selection rule of the form $\hat{\mu}(X)>\tilde{T}$ with $\tilde{T}$ determined by calibration scores, including the conformal $p$-value-based methods. We omit the details here for brevity, but one can verify that the $p$-value in (24) yields a selection rule of this form, and that the $p$-value in (23) is deterministically larger than (i.e., dominated by) the $p$-value in (24).